OpenStudy (anonymous):

simplify each of the following: √27 √162 √99 √486 5√8 -5√63

OpenStudy (anonymous):

\[\sqrt{27}=\sqrt{3}\sqrt{9}=3\sqrt{3}\]

OpenStudy (calculusfunctions):

@dr25161 please stop giving out answers!

OpenStudy (epikrebel):

\[\sqrt{162}=\sqrt{2*3*3*3*3}\]

OpenStudy (epikrebel):

\[\sqrt{486}=\sqrt{2*3*3*3*3*3}\]

OpenStudy (anonymous):

Aren't we supposed to give answers?

OpenStudy (calculusfunctions):

To simplify a square root, i). Find the greatest perfect square which the radicand is divisible by ii). write the square root as a product of the square roots of that perfect square and the remaining factor. iii). Simplify Example:\[\sqrt{288}\] Solution: i). The greatest perfect square which the radicand is divisible by is 144. ii). \[\sqrt{288}=\sqrt{144}\sqrt{2}\]iii).\[\sqrt{144}\sqrt{2}=12\sqrt{2}\]Therefore\[\sqrt{288}=12\sqrt{2}\]

OpenStudy (epikrebel):

This is open study true we give answers but we need to help them understand the material. This website isn't just about answers. Its about understanding the material.

OpenStudy (calculusfunctions):

@dr25161 NO, you're supposed to teach, explain with examples like I just demonstrated.

OpenStudy (anonymous):

Yeah, you're right. I'm just a bit lazy

OpenStudy (anonymous):

Thanks guys!! sooo much!!! this helped me heaps!!! :) :) :D

OpenStudy (anonymous):

Thanks you guys , i too had doubt in it :)